### v1

@@ Abstract
Divergence constraints are present in the governing equations of many physical phenomena, and they usually lead to a Poisson equation whose solution typically is the main bottleneck of many simulation codes. Algebraic Multigrid (AMG) is arguably the most powerful preconditioner for Poisson’s equation, and its effectiveness results from the complementary roles played by the smoother, responsible for damping high-frequency error components, and the coarse-grid correction, which in turn reduces low-frequency modes. This work presents several strategies to make AMG more compute-intensive by leveraging reflection, translational and rotational symmetries, often present in academic and industrial configurations. The best-performing method, AMGR, is based on a multigrid reduction framework that introduces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving AMG’s excellent convergence, AMGR allows replacing the standard sparse matrix-vector product with the more compute-intensive sparse matrix-matrix product, yielding significant accelerations. Numerical experiments on industrial CFD applications demonstrated up to 70% speed-ups when solving Poisson’s equation with AMGR instead of AMG. Additionally, strong and weak scalability analyses revealed no significant degradation.
@@ 1 Introduction
Divergence constraints are prevalent in physical problems, often following fundamental conservation principles like mass or electrical charge conservation. Such constraints lead to a Poisson equation that plays a fundamental role in many areas of science and engineering, such as computational fluid dynamics (CFD), linear elasticity, and electrostatics. Indeed, the solution of the associated linear system is generally the most computationally intensive part of scientific simulation codes, and the design and implementation of Poisson solvers are far from straightforward due to the interplay of numerical and hardware challenges.
@@
However, the sequential nature of such methods and the increasing availability of parallel computers in the late ’90s made them lose ground against other alternatives with higher degrees of parallelism. For instance, the application of preconditioners based on approximate inverses solely relies on the sparse matrix-vector product (SpMV), an easily parallelisable operation. In some cases, even their construction is reasonably concurrent. The most prominent variants are AINV [12, 13, 11], SPAI [25] and the Factored Sparse Approximate Inverse (FSAI) [33, 31]. While approximate inverses provide a high degree of parallelism, they are not optimal in the sense that when the mesh size decreases (hence increasing the linear system size), the problem becomes more ill-conditioned and more iterations are required to reach the same accuracy.
This problem worsens nowadays, as extreme-scale linear systems must be solved on massively parallel supercomputers, and single-level preconditioners such as incomplete factorisations or approximate inverses generally require excessive iterations. The problem of scalability is overcome with multilevel preconditioners like Geometric Multigrid (GMG) or Algebraic Multigrid (AMG) ][52, 55, 59]. Thanks to the interplay between smoother and coarse-grid correction, these methods often solve a given PDE with a number of iterations independent of the mesh size. Many freely available multigrid packages exist. To cite a few, Hypre [28], Trilinos [14], and PETSc [9] provide very effective implementations with excellent scalability.
@@
A few works exploiting symmetries for solving Poisson’s equation exist [5, 4, 24, 49]. This paper extends them by proposing enhanced variants of FSAI and AMG preconditioners. Their advantages result from mitigating several of the computational challenges above. Namely, we show that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) satisfy regular block structures. This, in turn, allows the standard SpMV to be replaced with the more compute-intensive sparse matrix-matrix product (SpMM[3], leading to lighter and faster variants of FSAI and AMG. Several strategies are considered, the best performing one based on a multigrid reduction framework that induces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving the excellent convergence of the standard AMG, replacing SpMV with SpMM yielded significant speed-ups.
@@
The remaining sections are organised as follows. Section 2 presents a strategy for exploiting reflection symmetries through low-rank corrections, and Section 3 extends it to translational and rotational symmetries. Section 4 discusses the parallel implementation, Section 5 presents meaningful numerical experiments, and Section 6 gives some concluding remarks.
@@ 2 Symmetry-aware spatial discretisation
Let us start by recalling a strategy to enhance FSAI and later extend it towards AMG. Such a strategy introduces low-rank corrections to enable the use of SpMM but requires a discretisation consistent with the spatial symmetries. Given an arbitrary mesh presenting a single reflection symmetry, we will order its grid points by first indexing the ones lying on one half and then those on the other. Therefore, analogous to Fig. 1, if we impose the same local ordering (mirrored by the symmetry’s hyperplane) to the resulting two subdomains, we ensure that all the scalar fields satisfy:
@@
(2) $H=\begin{pmatrix}H_{1,1}&H_{1,2}\\ H_{2,1}&H_{2,2}\end{pmatrix}\in\mathbb{R}^{n\times m},$
where $H_{i,j}\in\mathbb{R}^{n/2\times m/2}$ accounts for the couplings between the $i$th and $j$th subdomains. As long as $H$ only depends on geometric quantities (which is typically the case) or on material properties respecting the symmetries, given that both subdomains are identical and thanks to the mirrored ordering, we have that:
@@
where $H_{i}\in\mathbb{R}^{n/4\times m/4}$ contains the couplings between the first and $i$th subdomains.
Thanks to the discretisation presented above, exploiting $s$ reflection symmetries allows meshing a $1/2^{s}$ fraction of the entire domain, henceforth named base mesh. Figure 3 depicts a schematic representation of the base mesh on an arbitrary domain with two symmetries. Then, instead of building the entire operators, $H\in\mathbb{R}^{n\times m}$, it is only needed to build the base mesh’s couplings with itself, $H_{1}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$, and with its $12^{s}-$ 1 mirrored counterparts, $H_{2},\dots,H_{2^{s}}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$. As a result, both the setup and memory footprint of the matrices are reduced by a factor of $2^{s}$ [3].
@@
where $H_{\text{inn}}\coloneqq H_{1}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$ and $H_{\text{out}}\coloneqq H-\mathbb{I}_{2^{s}}\otimes H_{1}\in\mathbb{R}^{n\times m}$.
The splitting of Eq. 6 allows the standard SpMV by $H$ to be replaced with a specialised version of the more compute-intensive SpMM. The resulting kernel is a fusion of an SpMM by $H_{\text{inn}}$, an SpMV by $H_{\text{out}}$ and a linear combination of vectors (axpy), and has been employed throughout the simulations to make up to 5x faster their matrix multiplications [3]. In order to elucidate the computational advantages of SpMM, let us note that the standard approach for applying $\mathbb{I}_{2^{s}}\otimes H_{\text{inn}}$ to a vector is through an SpMV call, which performs the following operation:
@@ 2.1 Low-rank corrections for factorable preconditioners
The idea of applying low-rank corrections arises from the close similarity between $\hat{A}$’s subsystems. Indeed, in Eq. 14, all the outer-couplings are substantially sparser than the inner and, as discussed in [2], $\operatorname{rank}(A_{\text{out},i})=\mathcal{O}(n^{2/3})$, whereas $\operatorname{rank}(A_{\text{inn}})=\mathcal{O}(n)$. Hence, for all $i\in\{1,\dots,2^{s}\}$:
@@
It can be shown experimentally that rough and cost-effective approximations of $U_{k}$ and $\Lambda_{k}$ suffice for preconditioning purposes. Then, applying the above procedure to each of the $2^{s}$ subsystems leads to the following low-rank corrected FSAI, henceforth denoted as LRCFSAI($k$):
@@
Table 1 summarises the convergence of LRCFSAI($k$) on the model problem of Eq. 25. Clearly, low-rank corrections outweigh the fact of ignoring the outer-couplings, actually making LRCFSAI($k$) converge faster than the standard FSAI. As a result, we proceeded with its parallel implementation, which is out of the scope of this work but was thoroughly reviewed in [2].
@@ 2.2 Low-rank corrections for non-factorable preconditioners
Bearing in mind that the application of AMG relies on matrix multiplications, we aim to accelerate it by replacing SpMV with SpMM. However, AMG is not explicitly factorable, and we need to develop alternative low-rank corrections. With this aim, let us recall the following result [36].
@@
(27) $M_{\text{inn}}\simeq A_{\text{inn}}^{-1}.$
@@
(28) $Y\coloneqq\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}-M_{\text{inn}}\hat{A}_{i}\in\mathbb{R}^{n/2^{s}\times n/2^{s}},$
@@
(29) $\hat{A}_{i}^{-1}=\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U\left(\Sigma^{-1}-V^{T}U\right)^{-1}V^{T}\right)M_{\text{inn}}.$
@@
(31) $\hat{A}_{i}^{-1}\simeq\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U_{k}(\Sigma_{k}^{-1}-V_{k}^{T}U_{k})^{-1}V_{k}^{T}\right)M_{\text{inn}}=U_{k}\Theta_{k}V_{k}^{T}M_{\text{inn}},$
where we considered biorthonormal bases for the right- and left-eigenvectors, i.e., $V_{k}^{T}U_{k}=\mathbb{I}_{k}$, and defined $\Theta_{k}\coloneqq\mathbb{I}_{k}+(\Sigma_{k}^{-1}-\mathbb{I}_{k})^{-1}\in\mathbb{R}^{k\times k}$. Then, by recalling the Sherman-Morrison-Woodbury formula, we have that: $\Theta_{k}=\mathbb{I}_{k}+\Sigma_{k}+\Sigma_{k}(\mathbb{I}_{k}-\Sigma_{k})^{-1}\Sigma_{k}$.
When it comes to selecting the most relevant eigenvectors, let us note that, as in Section 2.1, $Y$ measures how far it is each preconditioned subsystem, $M_{\text{inn}}\hat{A}_{i}$, from the identity matrix. Then, given the harmful effect of small eigenvalues in the convergence of Krylov subspace methods, $Y$’s most effective eigenvectors are those associated with the smallest eigenvalues of $X\coloneqq M_{\text{inn}}\hat{A}_{i}$. Hence, by computing a low-rank approximation of $X$:
(32) $X\simeq U_{k}\Lambda_{k}V_{k}^{T},$
we can obtain the following truncated eigendecomposition of $Y$:
(33) $Y\simeq U_{k}(\mathbb{I}_{k}-\Lambda_{k})V_{k}^{T},$
which gives $\Sigma_{k}=\mathbb{I}_{k}-\Lambda_{k}$ and, consequently:
(34) $\Theta_{k}=\mathbb{I}_{k}+\left(\mathbb{I}_{k}-\Lambda_{k}\right)+\left(\mathbb{I}_{k}-\Lambda_{k}\right)\left(\mathbb{I}_{k}-\left(\mathbb{I}_{k}-\Lambda_{k}\right)\right)^{-1}\left(\mathbb{I}_{k}-\Lambda_{k}\right)=\Lambda_{k}^{-1},$
therefore matching the largest eigenvalues of the correction with the smallest eigenvalues of the preconditioned subsystem.
Ultimately, applying the above procedure to each of the $2^{s}$ subsystems results in the following low-rank corrected AMG, henceforth denoted as LRCAMG($k$):
(35) $\mathbb{I}_{2^{s}}\otimes M_{\text{inn}}+\begin{pmatrix}U_{k,1}\Theta_{k,1}V_{k,1}^{T}&&\\ &\ddots&\\ &&U_{k,2^{s}}\Theta_{k,2^{s}}V_{k,2^{s}}^{T}\end{pmatrix}.$
LRCAMG($k$) is compatible with SpMM in the application of $M_{\text{inn}}$. However, this comes at the price of introducing another level of approximation by ignoring $\hat{A}$’s outer-couplings. While low-rank corrections proved very effective on LRCFSAI($k$), even accelerating its convergence, (relatively) low-rank perturbations had a critical impact on AMG. Indeed, Table 2 summarises the results obtained with LRCAMG($k$) on the model problem of Eq. 25. Unfortunately, low-rank corrections cannot restore the effectiveness of AMG, which excels in removing error components on the lower part of the spectrum with high accuracy, something that low-rank correction cannot match. Due to its unsatisfactory performance, we abandoned its parallel implementation and explored the better alternatives of Section 3.

### v2

@@ Abstract
Divergence constraints are present in the governing equations of numerous physical phenomena, and they usually lead to a Poisson equation whose solution represents a bottleneck in many simulation codes. Algebraic Multigrid (AMG) is arguably the most powerful preconditioner for Poisson’s equation, and its effectiveness results from the complementary roles played by the smoother, responsible for damping high-frequency error components, and the coarse-grid correction, which in turn reduces low-frequency modes. This work presents several strategies to make AMG more compute-intensive by leveraging reflection, translational and rotational symmetries. AMGR, our final proposal, does not require boundary conditions to be symmetric, therefore applying to a broad range of academic and industrial configurations. It is based on a multigrid reduction framework that introduces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving AMG’s excellent convergence, AMGR allows replacing the standard sparse matrix-vector product with the more compute-intensive sparse matrix-matrix product, yielding significant accelerations. Numerical experiments on industrial CFD applications demonstrated up to 70% speed-ups when solving Poisson’s equation with AMGR instead of AMG. Additionally, strong and weak scalability analyses revealed no significant degradation.
@@ 1 Introduction
Divergence constraints are prevalent in physical problems, often following fundamental conservation principles like mass or electrical charge conservation. Such constraints lead to a Poisson equation that plays a fundamental role in many areas of science and engineering, such as computational fluid dynamics (CFD), linear elasticity, and electrostatics. Indeed, the solution of the associated linear system is usually the most computationally intensive part of scientific simulation codes, and the design and implementation of Poisson solvers are far from straightforward due to the interplay of numerical and hardware challenges.
@@
However, the sequential nature of such methods and the increasing availability of parallel computers in the late ’90s made them lose ground against alternatives with higher degrees of parallelism. For instance, the application of preconditioners based on approximate inverses solely relies on the sparse matrix-vector product (SpMV), an easily parallelisable operation. In some cases, even their construction is reasonably concurrent. The most prominent variants are AINV [11, 12, 10], SPAI [24] and the Factored Sparse Approximate Inverse (FSAI) [33, 30]. While approximate inverses provide a high degree of parallelism, they are not optimal in the sense that when augmenting the mesh resolution (hence increasing the linear system size), the problem becomes more ill-conditioned and more iterations are required to reach the same accuracy.
This problem worsens nowadays, as extreme-scale linear systems must be solved on massively parallel supercomputers, and single-level preconditioners such as incomplete factorisations or approximate inverses generally require excessive iterations. The problem of scalability is overcome with multilevel preconditioners like Geometric Multigrid (GMG) or Algebraic Multigrid (AMG) [52, 55, 59]. Thanks to the interplay between smoother and coarse-grid correction, these methods often solve a given PDE with a number of iterations independent of the mesh size. Many freely available multigrid packages exist. To cite a few, Hypre [27], Trilinos [13], and PETSc [8] provide very effective implementations with excellent scalability.
@@
This work presents multiple strategies leading to a multigrid reduction framework that mitigates several of the aforementioned computational challenges. Namely, we show that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) conform to regular block structures regardless of the boundary conditions. This, in turn, enables replacing the standard SpMV with the more compute-intensive sparse matrix-matrix product (SpMM).
Similar strategies were used in the context of CFD simulations [49], but they were limited to reflection symmetries, and their main goal was to reduce the memory footprint [3, 23]. In [4], such strategies were generalised by leveraging an arbitrary number of reflection symmetries to block diagonalise the discrete Laplacian, which improved iterative methods’ convergence and granted several computational advantages. Nevertheless, the resulting subsystems were not identical, preventing the use of SpMM on complex preconditioners. In [2], low-rank corrections were introduced to make FSAI compatible with SpMM, resulting in a much faster and significantly lighter variant. However, given the poor weak scaling of single-level preconditioners, the present work extends past strategies to multilevel preconditioning. Namely, to AMG, arguably the most powerful method for preconditioning Poisson’s equation.
Extending SpMM’s use to AMG is far from direct, and for completeness, intermediate unsatisfactory approaches leading to AMGR, our final proposal, are included in this work. AMGR relies on a multigrid reduction framework that introduces an aggressive coarsening to the top of the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother, which replaces SpMV with SpMM. Subsequent levels are purely algebraic. Apart from reflections, AMGR also leverages translational and rotational symmetries and, most notably, is compatible with problems with asymmetric boundary conditions, therefore applying to a significantly broader range of industrial applications, on which it proved up to 70% faster than the standard AMG thanks to preserving its excellent convergence while enjoying the computational advantages of SpMM.
@@
The remaining sections are organised as follows. Section 2 recalls a strategy for exploiting reflection symmetries through low-rank corrections, and extends it to AMG. Then, Section 3 introduces an ordering able to cope with asymmetric boundary conditions, as well as translational and rotational symmetries, and ultimately leading to AMGR. Section 4 discusses the parallel implementation, Section 5 presents numerical experiments on industrial CFD applications, and Section 6 gives some concluding remarks.
@@ 2 Symmetry-aware spatial discretisation
Let us start by recalling a strategy to enhance FSAI [2] and later extend it towards AMG. Such a strategy introduces low-rank corrections to enable the use of SpMM but requires a discretisation consistent with the spatial symmetries. Although boundary conditions need not be symmetric, they must be of the same type. Nevertheless, the final approach of Section 3.2 will free this constraint, too. Then, given an arbitrary mesh presenting a single reflection symmetry, we will order its grid points by first indexing the ones lying on one half and then those on the other. Therefore, analogous to Fig. 1, if we impose the same local ordering (mirrored by the symmetry’s hyperplane) to the resulting two subdomains, we ensure that all the scalar fields satisfy:
@@
(2) $H=\begin{pmatrix}H_{1,1}&H_{1,2}\\ H_{2,1}&H_{2,2}\end{pmatrix}\in\mathbb{R}^{n\times n},$
where $H_{i,j}\in\mathbb{R}^{n/2\times n/2}$ accounts for the couplings between the $i$th and $j$th subdomains. As long as $H$ only depends on geometric quantities (which is typically the case) or on material properties respecting the symmetries, given that both subdomains are identical and thanks to the mirrored ordering, we have that:
@@
where $H_{i}\in\mathbb{R}^{n/4\times n/4}$ contains the couplings between the first and $i$th subdomains.
Thanks to the discretisation presented above, exploiting $s$ reflection symmetries allows meshing a $1/2^{s}$ fraction of the entire domain, henceforth named base mesh. Figure 3 depicts a schematic representation of the base mesh on an arbitrary domain with two symmetries. Then, instead of building the entire operators, $H\in\mathbb{R}^{n\times n}$, it is only needed to build the base mesh’s couplings with itself, $H_{1}\in\mathbb{R}^{n/2^{s}\times n/2^{s}}$, and with its $12^{s}-$ 1 mirrored counterparts, $H_{2},\dots,H_{2^{s}}\in\mathbb{R}^{n/2^{s}\times n/2^{s}}$. As a result, both the setup and memory footprint of the matrices are reduced by a factor of $2^{s}$ [3].
@@
where $H_{\text{inn}}\coloneqq H_{1}\in\mathbb{R}^{n/2^{s}\times n/2^{s}}$ and $H_{\text{out}}\coloneqq H-\mathbb{I}_{2^{s}}\otimes H_{1}\in\mathbb{R}^{n\times n}$.
The splitting of Eq. 6 allows the standard SpMV by $H$ to be replaced with a specialised version of the more compute-intensive SpMM. The resulting kernel is a fusion of an SpMM by $H_{\text{inn}}$, an SpMV by $H_{\text{out}}$ and a linear combination of vectors, and has been employed throughout the simulations to make up to 5x faster their matrix multiplications [3]. In order to elucidate the computational advantages of SpMM, let us note that the standard approach for applying $\mathbb{I}_{2^{s}}\otimes H_{\text{inn}}$ to a vector is through an SpMV call, which performs the following operation:
@@ 2.1 Low-rank corrections for factorable preconditioners
The idea of applying low-rank corrections arises from the close similarity between $\hat{A}$’s subsystems. Indeed, in Eq. 14, all the outer-couplings are substantially sparser than the inner and, as discussed in [2], assuming a $d$-dimensional domain with a similar scale in all directions, we have that $\operatorname{rank}(A_{\text{out},i})=\mathcal{O}(n^{(d-1)/d})$, whereas $\operatorname{rank}(A_{\text{inn}})=\mathcal{O}(n)$. Hence, for all $i\in\{1,\dots,2^{s}\}$:
@@
It can be shown experimentally that rough and cost-effective approximations of $U_{k}$ and $\Lambda_{k}$ suffice for preconditioning purposes. Then, applying the above procedure to each of the $2^{s}$ subsystems leads to the following low-rank corrected FSAI, henceforth denoted as Low-Rank Corrected FSAI (LRCFSAI($k$)):
@@
Table 1 shows the convergence of LRCFSAI($k$) on the model problem of Eq. 25. The results correspond to a MATLAB implementation using an FSAI with up to 50 nonzeros per row and a pre-filtering tolerance $\tau=0.01$ [30]. Clearly, low-rank corrections outweigh the fact of ignoring the outer-couplings, actually making LRCFSAI($k$) converge faster than the standard FSAI. As a result, we proceeded with its parallel implementation, which is out of the scope of this work but was thoroughly reviewed in [2].
@@ 2.2 Low-rank corrections for non-factorable preconditioners
Bearing in mind that the application of AMG relies on matrix multiplications, we aim to accelerate it by replacing SpMV with SpMM. However, AMG is not explicitly factorable, and we need to develop alternative low-rank corrections. With this aim, let us recall the following result [36], in which we replace SVD with eigendecomposition, as it facilitated the selection of the correction vectors.
@@
(27) $M_{\text{inn}}^{-1}\simeq A_{\text{inn}}^{-1}.$
@@
(28) $Y\coloneqq\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}-M_{\text{inn}}^{-1}\hat{A}_{i}\in\mathbb{R}^{n/2^{s}\times n/2^{s}},$
@@
(29) $\hat{A}_{i}^{-1}=\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U\left(\Sigma^{-1}-V^{T}U\right)^{-1}V^{T}\right)M_{\text{inn}}^{-1}.$
@@
(31) $\hat{A}_{i}^{-1}\simeq\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U_{k}(\Sigma_{k}^{-1}-V_{k}^{T}U_{k})^{-1}V_{k}^{T}\right)M_{\text{inn}}^{-1}=\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U_{k}\Theta_{k}V_{k}^{T}\right)M_{\text{inn}}^{-1},$
where we considered biorthonormal bases for the right- and left-eigenvectors, i.e., $V_{k}^{T}U_{k}=\mathbb{I}_{k}$, and defined $\Theta_{k}\coloneqq(\Sigma_{k}^{-1}-\mathbb{I}_{k})^{-1}\in\mathbb{R}^{k\times k}$.
Then, applying the above procedure to each of the $2^{s}$ subsystems results in the following low-rank corrected AMG, henceforth denoted as Low-Rank Corrected AMG (LRCAMG($k$)):
(32) $\left[\mathbb{I}_{n}+\begin{pmatrix}U_{k,1}\Theta_{k,1}V_{k,1}^{T}&&\\ &\ddots&\\ &&U_{k,2^{s}}\Theta_{k,2^{s}}V_{k,2^{s}}^{T}\end{pmatrix}\right]\left(\mathbb{I}_{2^{s}}\otimes M_{\text{inn}}^{-1}\right).$
When it comes to selecting the most relevant eigenvectors, let us note that, as in Section 2.1, $Y$ measures how far it is each preconditioned subsystem, $M_{\text{inn}}^{-1}\hat{A}_{i}$, from the identity matrix. Then, given the harmful effect of small eigenvalues in the convergence of Krylov subspace methods, $Y$’s most effective eigenvectors are those associated with the smallest eigenvalues of $X\coloneqq M_{\text{inn}}^{-1}\hat{A}_{i}$. Additionally, by computing the eigendecomposition of $X$:
(33) $X=U\Lambda V^{T},$
we can obtain the following eigendecomposition of $Y$:
(34) $Y=U(\mathbb{I}-\Lambda)V^{T},$
which gives $\Sigma=\mathbb{I}-\Lambda$. Hence, the smallest, and harmful, eigenvalues of $X$ correspond to the largest positive eigenvalues of $Y$, whereas large (in magnitude) negative eigenvalues of $Y$ correspond to the largest, and harmless, eigenvalues of $X$.
LRCAMG($k$) is compatible with SpMM in the application of $M_{\text{inn}}^{-1}$. However, this comes at the price of introducing another level of approximation by ignoring $\hat{A}$’s outer-couplings. While low-rank corrections proved very effective on LRCFSAI($k$), even accelerating its convergence, (relatively) low-rank perturbations had a critical impact on AMG. Indeed, Table 2 summarises the results obtained with LRCAMG($k$) on the model problem of Eq. 25. The results correspond to a MATLAB implementation using an AMG with a PMIS coarsening [18], Extended+I interpolation [17] and an FSAI smoother with 15 nonzeros per row. Unfortunately, low-rank corrections cannot restore the effectiveness of AMG, which excels in removing error components on the lower part of the spectrum with high accuracy, something that low-rank correction cannot match. Due to its unsatisfactory performance, we abandoned its parallel implementation and explored the better alternatives of Section 3.